Let A and B be any two matrices such that B is the inverse of A.
3.1 Determine the relationship between the adjoint of A and the adjoint of B. (5)
3.2 Determine the relationship between the transpose of A and the transpose of B.
1
Expert's answer
2020-04-23T19:06:05-0400

Given that , "A^{-1}=B" .

1) We known that,

"\text{adj}(PQ)=(\text{adj}Q)(\text{adj}P) , \text{for any square matrix } P,Q \ \text{of the same order}."

And "\text{adj}(I)=I" ,where "I" is the identity matrix.

Since "AB=I=BA"

Therefore, "\text{adj}(AB)=adj(I)=adj(BA)."

"\implies (adjB)(adjA)=I=(adjA)(adjB)"

"\implies (adjB)=(adjA)^{-1}" .


2) We known that "(PQ)^T=Q^TP^T" for any square matrix "P,Q" of same order .

Since ,"AB=I=BA"

"\implies (AB)^T=I^T=(BA)^T"

"\implies B^TA^T=I=A^TB^T," as "I^T=I"

"\implies B^T={(A^T)}.^{-1}"